OpenStudy (anonymous):

Y% of 450 please help

jhonyy9 (jhonyy9):

how many is y ?

OpenStudy (anonymous):

i dont know, it says to write it as an expression @jhonyy9

jhonyy9 (jhonyy9):

so ok than you need to us formula the rule of 3 simple 450 ------ 100% x --------- Y % ------------------ x = (y*450)/100 = Y*4,5 so here x mean the Y% of 450 hope that is understandably sure easy

OpenStudy (anonymous):

thank you @jhonyy9

jhonyy9 (jhonyy9):

was my pleasure bye